Good day
can someone help me with this function
g(x)= x^2 + x -7/3
Which is the domain?

Answer by ewatrrr(10682)


 
Hi,
g(x)= x^2 + x -7/3 = (x+1/2)^2 - 1/4 - 7/3 =(x+1/2)^2 - 31/12
domain: all real numbers.
range: g(x) ≥ - 31/12